# D4C: Improving Negative Example Quality to Enhance Machine Abstract Reasoning Ability

### v14

@@ V-A3 Concept Consistency Measurement Module
The inference module is capable of transforming incomplete representation matrices into sampled results of Gaussian Mixture Models (GMMs) with five components, treating these samples as the fuzzy representations of inherent concepts within the incomplete matrices. A complete $3\times 3$ matrix can give rise to nine distinct incomplete matrices; hence, we can extract nine sets of concepts from a single complete matrix, obtaining their fuzzy representations. Given that these fuzzy representations of concepts follow a mixture of Gaussian distributions, the task of the Concept Consistency Measurement Module is to calculate the distances between these GMMs based on the Sinkhorn distance [39], and to use this distance to reflect the consistency score of the concepts within a complete $3\times 3$ matrix.
@@ V-A5 Preliminary experiments in Lico-Net
After clarifying the structure of Lico-Net and its training loss function, we conducted preliminary experiments across all sub-datasets of RAVEN. Rather than focusing solely on the inference accuracy of Lico-Net, we placed greater emphasis on how Lico-Net learns the embedded concepts within the RAVEN dataset. Since Lico-Net can extract nine sets of concepts from a single complete $3\times 3$ matrix, it is worth delving into its concept extractor’s performance across the entire dataset.
Therefore, we organized this special experiment. Taking the “Distribution 9” configuration in RAVEN as an example, 20 problem matrices with identical rules were generated for this configuration. Subsequently, we assessed the distances between concepts within and between these 20 problem matrices. As shown in Figure 8, despite the identical rules for these 20 problem matrices, it is observed from Lico-Net’s perspective that only the concepts within the same matrix are consistently matched. In other words, Lico-Net’s induction of concepts or patterns is far from human-like. Whether this pattern mismatch is a negative effect brought by concept fuzzification is worth discussing.
@@ V-B A novel baseline, Lico-Net-Bongard, for Bongard-Logo
The designers of Bongard-logo drew inspiration from high-dimensional human concepts and knowledge, carefully categorizing the Bongard-logo database into three distinct problem types: FF, BA, and HD. It is noteworthy that during the testing process, the HD concept is further decomposed into NV and CM. Bongard-Logo requires participants to analyze the clustering logic between the primary and auxiliary sets outlined in the problem statement, and to correctly categorize two given images into these sets. For this type of problem, our approach, which involves extracting concepts from images as fuzzy representations, is likely to be effectively applied. By utilizing a deep neural network to extract and represent concepts from images as fuzzy representations, we can naturally impose constraints on the distances between these concepts to align with the problem-solving logic and clustering logic of Bongard-Logo. This ensures that concepts within the primary set are more closely related, while concepts within the auxiliary set are distinct from those in the primary set.
Based on our methodology, we have designed Lico-Net-Bongard.
@@ V-B1 The structure of Lico-Net-Bongard
For a clearer exposition of the Lico-Net-Bongard architecture, Figure 9 has been constructed. In Figure 9 (a), we detail the annotation of the 14 images within a single Bongard-Logo problem, which is instrumental in defining the loss function for training Lico-Net-Bongard. Figure 9 (b) illustrates the process by which Lico-Net-Bongard extracts concepts from individual images in Bongard-Logo. The process is as follows: after utilizing ResNet18 to convert Bongard-Logo images ($x_{i}$) into feature maps ($h_{i}$), Lico-Net-Bongard scatters these feature maps along the receptive field into vector groups ($h_{ij}$ where $j$ denotes the receptive field index and $m$ represents the size of the receptive field), which are then processed by a standard Transformer-Encoder to obtain the multi-viewpoint representations ($h^{\prime}_{ij}$) of the image. Thereafter, using a fuzzification head, these multi-viewpoint representations are individually transformed into distributions ($z_{ij}$). Additionally, reparameterization techniques are employed to compute the fuzzy representations of the concepts from the transformed distributions.
@@ V-B2 Loss function for Lico-Net-Bongard
After Lico-Net-Bongard processes all 14 images ($\{x_{i}|\,i\in[1,14]\}$) in a Bongard-Logo case to form concepts ($\{z_{ij}|\,i\in[1,14],j\in[1,m]\}$), we can optimize Lico-Net-Bongard by constraining the distances between these concepts. The specific loss function design can be expressed as:
@@
The expression for $\ell_{\mathbf{Lico}}$ is straightforward: it calculates the distances between each pair of concepts within the set $\{z_{ij}\}$ and constrains these distances using cross-entropy loss. Here, $D(\alpha,\beta)$ represents the function that calculates the Sinkhorn distance between two distributions $\alpha$ and $\beta$. Similarly, we employ the reparameterization method to sample from $z_{ij}$, with the number of samples set to 10.
@@ V-B3 Preliminary experiments in Lico-Net-Bongard
With the structure of Lico-Net-Bongard and its training loss function clearly defined, we conducted preliminary experiments on the Bongard-Logo dataset. We carried out two sets of experiments, one where the four types of concepts from the Bongard-Logo database were mixed together, and another where they were separated, to observe the performance of Lico-Net-Bongard. The experimental results are presented in Table I.
@@
The NICO database presents a classic visual discrimination task, but the data within it exhibit a shift between the distribution of background pixels and the distribution of foreground subjects. In NICO, the “domain” refers to the background element, which can undergo two types of shifts: Diversity Shift and Correlation Shift. Diversity Shift occurs when the test set contains new, unseen background styles. For example, a model trained to recognize sheep may struggle if it encounters sheep in snow, a scenario not present in the training data. Correlation Shift, on the other hand, involves the relationship between background and foreground pixel distributions in the training set. If most sheep images in the training data have a grassland background, a model might mistakenly associate the label with the background rather than the foreground, leading to impaired performance.
@@ V-C1 Preliminary experiments in NICO
Diversity Shift: We randomly select 8 domains as source domains and the remaining 2 as target domains. After training and testing basic deep learning models on all 10 domains to establish a benchmark, we train the model on source domains and test it on target domains. By comparing the results to the benchmark, we observe the impact of Diversity Shift. For a fair comparison, we maintain consistent data volumes in all training and testing sets. The experimental results in Table II highlight the significant effect of Diversity Shift on deep learning models.
@@
Correlation Shift: In our experiment, we exacerbate this shift by randomly downsampling the training set. Specifically, we randomly select one of the 8 source domains as the primary and reduce the sample sizes of the remaining 7 to 1/5 of the primary. We then train on the reduced source domains and test on the unchanged target domain. Table III shows Vgg16 results on the NICO dataset with more pronounced correlation shift.
@@ V-D Discussion on Preliminary Experiments
In the Bongard-Logo problem, different concepts require Lico-Net-Bongard to store different image pixel encoding patterns. When these encoding patterns are mutually exclusive, the mixed concepts confuse Lico-Net-Bongard, making it uncertain which concept to rely on when solving problems during testing. This transforms four concepts that are clear and organized from a human perspective into four conflicting concepts from the perspective of Lico-Net-Bongard. As a result, the phenomenon of concept mismatch observed in the concept dissection experiment becomes understandable. However, the design of the correct answers in the RAVEN problem is indeed based on its inherent rules and problem-solving logic, and the annotations in NICO are also accurately aligned with the foreground subjects. Therefore, the question arises: why do both RAVEN and NICO exhibit a degree of conceptual mismatch?
@@ VI Attribution Discussion of Mismatch in Deep-level Concepts (D2C)
Our preliminary experiments have revealed a significant finding: there exists a mismatch between the way our designed inference engine induces deep or implicit concepts in the data and the way humans interpret these concepts. This mismatch is one of the key factors limiting the performance of the inference engine. Therefore, delving into the causes of the concept mismatch and devising corresponding solutions will be an effective approach to enhancing the performance of the inference engine.
Upon analysis, one of the reasons for concept mismatch is attributed to the low quality of the negative examples set in the problems.
@@ VI-1 RPM problem
Taking the RPM problem, most end-to-end RPM inference engine, including Lico-Net, follow a pattern: when the correct option is filled into the problem matrix, the inference engine returns a high probability value for this matrix, and conversely, returns a low probability value. This indicates that the essence of an RPM inference engine is to perform a conditional probability distribution over the options, conditioned on the problem matrix. The optimization of the inference engine can be seen as an adjustment of this conditional probability distribution. If we assume that the form of this conditional probability distribution, after determining the problem matrix, is a Gaussian distribution, then the optimization of this distribution can be regarded as adjusting the mean $\mu_{\theta}$ and variance $\sigma_{\theta}^{2}$ in the following formula:
@@
It can be observed that when positive samples (correct option) are substituted into the formula 2, we aim for the smallest possible result, which corresponds to minimizing the Euclidean distance between option $x_{\text{option}}$ and the mean of the correct solution $\mu_{\theta}$. Conversely, when negative samples (incorrect options) are substituted, we desire the largest possible result, equivalent to minimizing the standard deviation $\sigma_{\theta}^{2}$. It is well understood that the mean $\mu_{\theta}$ describes the center of a Gaussian distribution, while variance $\sigma_{\theta}^{2}$ describes the spread or dispersion. Consequently, it is not difficult to comprehend that the learning objective of end-to-end RPM inference engine like Lico-Net is a probability distribution delimited by positive and negative examples. In other words, the learning goal of these scoring-based end-to-end inference engine is not the Inference logic established by humans, but rather the distinction between positive and negative examples.
This indicates that when the quality of negative examples is not sufficiently high or not deceptive enough, it can lead to the engine learning concepts with relatively loose boundaries, which also causes the learned concepts to deviate significantly from the reasoning logic established by humans.
@@ VI-2 Bongard-Logo problem
This issue has also impacted the process of solving and learning the Bongard-Logo problem. By examining Equation 1, it is evident that the primary learning focus of Lico-Net-Bongard remains on ensuring that the similarity among concepts extracted from positive examples is higher than that between positive and negative examples. In other words, its learning objective is to distinguish the differences between positive and negative examples, rather than directly grasping the four specific concepts defined by humans in Bongard-Logo. Consequently, due to the inadequate quality of the negative examples provided by Bongard-Logo, it communicates a set of conflicting concepts.
@@ VI-4 Summary
In summary, most end-to-end deep learning models focus more on the differences between positive and negative examples in the data, rather than the reasoning logic behind why positive examples are correct and negative examples are incorrect. When the quality of negative examples in the data is not high enough, the distinction between positive and negative examples is unable to accurately convey the underlying logic.
@@ VII-A RPM and Bongard-Logo
When addressing the RPM problem and the Bongard-Logo problem, for a given observed sample, we will select images from two other distinct observed samples and incorporate them into the option pool of the original sample. The detailed procedure is illustrated in Figure 10.
@@
In Figure 9, (a) illustrates the approach for handling RPM problems, while (b) depicts the method for Bongard-Logo problems. Due to computational constraints, this paper opts to augment the negative examples with three additional samples for RPM and two additional samples for Bongard-Logo.
@@ VII-B NICO
For NICO, we cannot introduce images from outside the original dataset as negative samples, as this would disrupt the distribution of the original NICO data and undermine its significance. This is also the reason why we adhere to not importing pre-trained parameters for the deep learning model. However, our approach of introducing higher-quality negative samples remains unchanged.
When processing NICO data with a basic deep learning model, the comparison between positive and negative samples is implicitly achieved through cross-entropy. Therefore, this paper proposes that, when deep learning models process NICO data, we should introduce an explicit process of positive and negative sample comparison for the models. This is a way to improve the quality of negative samples without introducing additional data or extra supervisory signals. The explicit positive and negative sample comparison method proposed in this paper is the The Group-based Step-wise Comparison Learning Method (GSCLM). The design of GSCLM is quite straightforward. It separates the training of the perceptual part (convolutional layers) and the inductive part (fully connected layers) of the visual deep learning model, such as Vgg16.
@@ VII-B1 The training of the perceptual part
In the figure, “$\phi$” represents the perceptual part of a visual deep learning network.
Subsequently, we impose loss constraints on this representation group to optimize the perceptual component of the deep network. The loss constraints can be expressed as:
@@ VIII Designing More Demanding Negative Examples to Drive Deep-Level Concept Learning (D4C)
We have analyzed that the reason why data fails to convey precise concepts for deep models to learn is the poor quality of negative examples. So how can we obtain higher-quality negative examples without introducing additional new data or more supervisory signals? We have emphasized the importance of the quality of negative examples multiple times, and in fact, the metric reflecting the quality of an example is the probability feedback provided by the inference engine. Therefore, this paper proposes that we can design an additional sample generator to produce negative examples that receive higher evaluations from the inference engine compared to the original negative examples. The framework that integrates the generator with the inference engine can be illustrated by Figure 12.
Figure 12: The framework that integrates the generator with the inference engine
@@
Formulas (5) and (6) implicitly involve the participation of $k_{n-1}$ and $V_{n-1}$. This paper use element-wise ELU + 1 as the activation function $\sigma(\cdot)$. After the step-wise linear attention mechanism calculation is completed, the resulting $O_{s}$ is processed by a Feedforward layer, which follows the classic design from the Transformer, and then summed with $Q^{\prime}_{n}$ to obtain the updated promoter. The updated promoter can serve as the representation of the generated negative example, or it can also act as the input promoter for the block stacked after the current block.
@@
The stacked negative example generator is trained in tandem with our inference engine, Lico-Net or Lico-Net-Bongard, to improve their inference accuracy. During training, the inference engine and the negative example generator are trained alternately, which is consistent with the standard adversarial learning approach. It is noteworthy that during the generator’s training, we aspire for the generated negative examples to score higher in the inference engine than the original negative examples. Therefore, we utilize cross-entropy loss to optimize the generator. Based on the principle that unsupervised learning does not ensure disentanglement [45], the design of such a loss function indicates that this paper is not concerned with the specific attributes or concepts conveyed by the generated negative examples. Instead, it leverages the fuzziness of these negative examples to push the boundaries of concepts, which constitutes a contribution to the methodology of fuzzy systems.

### v16

@@ V-A3 Concept Consistency Measurement Module
The Reasoning Module and the Concept Extraction Module are capable of transforming incomplete representation matrices into concepts represented by five Gaussian distributions. Samples drawn using the reparameterization trick[7] from these five distributions are regarded as the probabilistic representations of these concepts. A complete $3\times 3$ matrix can generate nine distinct incomplete matrices; hence, we can extract nine sets of distribution-form concepts from a complete $3\times 3$ matrix and obtain the probabilistic representations for each set of concepts. The task of the Concept Consistency Measurement Module is to calculate the Sinkhorn distances [39] among these nine distribution-form concepts, based on their probabilistic representations, and to utilize this distance to reflect the consistency score of the concepts within a complete $3\times 3$ matrix. In this paper, we extract 10 probabilistic representations from each distribution-form concept to support the calculation of the Sinkhorn distance.
@@ V-A5 Preliminary experiments in Lico-Net
After clarifying the structure of Lico-Net and its training loss function, we conducted preliminary experiments across all sub-datasets of RAVEN. Rather than focusing solely on the reasoning accuracy of Lico-Net, we placed greater emphasis on how Lico-Net learns the embedded concepts within the RAVEN dataset. Since Lico-Net can extract nine sets of concepts from a single complete $3\times 3$ matrix, it is worth delving into its concept extractor’s performance across the entire dataset. Therefore, we organized this special experiment. Taking the “Distribution 9” configuration in RAVEN as an example, 20 problem matrices with identical rules were generated for this configuration. Subsequently, we assessed the distances between concepts within and between these 20 problem matrices. As shown in Figure 8, despite the identical rules for these 20 problem matrices, it is observed from Lico-Net’s perspective that only the concepts within the same matrix are consistently matched. In other words, Lico-Net’s induction of concepts or patterns is far from human-like. Whether this pattern mismatch is a negative effect brought by probabilistic representation learning is worth discussing.
@@ V-B A novel baseline, Lico-Net-Bongard, for Bongard-Logo
The designers of Bongard-logo drew inspiration from high-dimensional human concepts, categorizing the database into three problem types: FF, BA, and HD. (The HD concept will be split into the CM and NV concepts during testing. In other words, when training the reasoning engine, the three concepts of FF, BA, and HD are involved, while during testing, the performance of the reasoning engine will be evaluated on four concepts of FF, BA, CM, and NV.) Bongard-Logo tasks participants with analyzing clustering logic between primary and auxiliary sets and categorizing images accordingly.
We have designed a reasoning engine based on a deep neural network named Lico-Net-Bongard, which can extract abstract concepts in Bongard-Logo images and represent them as probabilistic representations. To ensure that these concepts are highly aligned with the problem-solving logic and clustering logic of Bongard-Logo, we further constrain the distances between the extracted concepts. The imposition of these constraints aims to ensure that the extracted concepts within the primary set are more closely connected, while clearly distinguishing the concepts in the auxiliary set from those in the primary set.
@@ V-B1 The structure of Lico-Net-Bongard
For a clearer exposition of the Lico-Net-Bongard architecture, Figure 9 has been constructed.
  1. 1.
    In Figure 9 (a), we detail the annotation of the 14 images within a single Bongard-Logo problem, which is instrumental in defining the loss function for training Lico-Net-Bongard.
  2. 2.
    Figure 9 (b) illustrates the process by which Lico-Net-Bongard extracts distribution-form concepts from individual images in Bongard-Logo. The process is as follows: after utilizing ResNet18 to convert Bongard-Logo images ($x_{i}$) into feature maps ($h_{i}$), Lico-Net-Bongard scatters these feature maps along the receptive field into vector groups ($h_{ij}$ where $j$ denotes the receptive field index and $m$ represents the size of the receptive field), which are then processed by a standard Transformer-Encoder to obtain the multi-viewpoint representations ($h^{\prime}_{ij}$) of the images ($x_{i}$). Thereafter, using a probabilistic head, these multi-viewpoint representations are individually transformed into distributions ($z_{ij}$). The samples drawn from these distributions using reparameterization techniques are regarded as the probabilistic representations of the concepts embedded in the Bongard-Logo images.
@@ V-B2 Loss function for Lico-Net-Bongard
After Lico-Net-Bongard processes all 14 images ($\{x_{i}|\,i\in[1,14]\}$) in a Bongard-Logo case to form concepts ($\{z_{ij}|\,i\in[1,14],j\in[1,m]\}$), we can optimize Lico-Net-Bongard by constraining the Sinkhorn distances between these distribution-form concepts. The specific loss function design can be expressed as:
@@
The expression for $\ell_{\mathbf{Lico}}$ is straightforward: it calculates the distances between each pair of concepts within the set $\{z_{ij}\}_{i=1}^{7}$ and constrains these distances using cross-entropy loss according to the problem-solving logic. Here, $D(\alpha,\beta)$ represents the function that calculates the Sinkhorn distance between two distributions $\alpha$ and $\beta$. Similarly, calculating the Sinkhorn distance requires obtaining observed samples from the distribution, which still involves using the reparameterization trick to sample several probabilistic representations from the distribution-form concepts $z_{ij}$, with the number of samples set to 10.
@@ V-B3 Preliminary experiments in Lico-Net-Bongard
With the structure of Lico-Net-Bongard and its training loss function clearly defined, we conducted preliminary experiments on the Bongard-Logo dataset. We carried out two sets of experiments, one where the four types of concepts from the Bongard-Logo database were mixed together, and another where they were separated, to observe the performance of Lico-Net-Bongard. The experimental results are shown in Table I.
@@
The NICO database[18] presents a classic visual discrimination task, but it exhibits a distribution shift between background and foreground pixels. In NICO, “domain” refers to the background, which can undergo two types of shifts: Diversity Shift, where the test set contains background styles not seen in the training set; and Correlation Shift, where a predominant association between background and foreground in the training data leads the visual model to incorrectly associate labels with the background, resulting in decreased performance.
@@ V-C1 Preliminary experiments in NICO
Diversity Shift: We randomly select 8 domains as the source domains, with the remaining 2 domains serving as the target domains. We first train and test some basic deep learning models on all 10 domains, using the test accuracies as a benchmark. Then, we train the model with data from the source domains and test it on the target domains, comparing the test results with the benchmark to observe the impact of Diversity Shift. To ensure a fair comparison, we maintain consistent data quantities in the training and testing sets across all experimental setups. The experimental results shown in Table II indicate the significant impact of Diversity Shift on deep learning models.
@@
Correlation Shift: In the experiment investigating the impact of Correlation Shift, we amplify this phenomenon by randomly reducing samples in some training sets. Specifically, we select one of the 8 source domains as the primary domain and reduce the number of samples in the remaining 7 domains to $1/5$ of the primary domain, making them secondary domains. We use the reduced source domains (primary and secondary) for training, while the target domain remains unchanged. The experimental results of Vgg16 on NICO with a stronger domain shift are shown in Table III, which we use as a benchmark for the Correlation Shift problem.
@@ V-D Discussion on Preliminary Experiments
In the Bongard-Logo problem, Lico-Net-Bongard may need to store different pixel encoding patterns for different concepts. When these patterns are mutually exclusive, mixed concepts can confuse the model, leading to uncertainty in problem-solving. This transforms four concepts that are clear and organized from a human perspective into four conflicting concepts from the perspective of Lico-Net-Bongard. As a result, the phenomenon of concept mismatch observed in the concept dissection experiment becomes understandable. However, unlike the Bongard-Logo problem, the correct answers for different instances in the RAVEN problem are all designed based on the same set of inherent rules and problem-solving logic, and the annotations in NICO are also accurately aligned with the foreground subjects. Thus, the question arises: why do RAVEN and NICO also exhibit a certain degree of concept mismatch?
@@ VI Attribution Discussion of Mismatch in Deep-level Concepts (D2C)
Our preliminary experiments have revealed a significant finding: there exists a mismatch between the way our designed reasoning engine induces deep or implicit concepts in the data and the way humans interpret these concepts. This mismatch is one of the key factors limiting the performance of the reasoning engine. Therefore, delving into the causes of the concept mismatch and devising corresponding solutions will be an effective approach to enhancing the performance of the reasoning engine. Upon analysis, one of the reasons for concept mismatch is attributed to the low quality of the negative examples set in the problems.
@@ VI-1 RPM problem
Taking the RPM problem, most end-to-end RPM reasoning engine, including Lico-Net, follow a pattern: when the correct option is filled into the problem matrix, the reasoning engine returns a high probability value for this matrix, and conversely, returns a low probability value. This indicates that the essence of an RPM reasoning engine is to perform a conditional probability distribution over the options, conditioned on the problem matrix. The optimization of the reasoning engine can be seen as an adjustment of this conditional probability distribution. If we assume that the form of this conditional probability distribution, after determining the problem matrix, is a Gaussian distribution, then the optimization of this distribution can be regarded as adjusting the mean $\mu_{\theta}$ and variance $\sigma_{\theta}^{2}$ in the following formula:
@@
It can be observed that when positive samples (correct option) are substituted into the formula 2, we aim for the smallest possible result, which corresponds to minimizing the Euclidean distance between option $x_{\text{option}}$ and the mean of the correct solution $\mu_{\theta}$. Conversely, when negative samples (incorrect options) are substituted, we desire the largest possible result, equivalent to minimizing the standard deviation $\sigma_{\theta}^{2}$. It is well understood that the mean $\mu_{\theta}$ describes the center of a Gaussian distribution, while variance $\sigma_{\theta}^{2}$ describes the spread or dispersion. Consequently, it is not difficult to comprehend that the learning objective of end-to-end RPM reasoning engine like Lico-Net is a probability distribution delimited by positive and negative examples. In other words, the learning goal of these scoring-based end-to-end reasoning engine is not the reasoning logic established by humans, but rather the distinction between positive and negative examples. This indicates that when the quality of negative examples is not sufficiently high or not deceptive enough, it can lead to the engine learning concepts with relatively loose boundaries, which also causes the learned concepts to deviate significantly from the reasoning logic established by humans.
@@ VI-2 Bongard-Logo problem
This issue has also impacted the process of solving and learning the Bongard-Logo problem. By examining Equation (1), it is evident that the primary learning focus of Lico-Net-Bongard remains on ensuring that the similarity among concepts extracted from positive examples is higher than that between positive and negative examples. In other words, its learning objective is to distinguish the differences between positive and negative examples, rather than directly grasping the four specific concepts defined by humans in Bongard-Logo. Consequently, due to the inadequate quality of the negative examples provided by Bongard-Logo, it communicates a set of conflicting concepts.
@@ VI-4 Summary
In summary, most end-to-end deep learning models focus more on the differences between positive and negative examples in the data, rather than the reasoning logic behind why positive examples are correct and negative examples are incorrect. When the quality of negative examples in the data is not high enough, the distinction between positive and negative examples is unable to accurately convey the underlying logic and embedded concepts.
@@ VII-A RPM and Bongard-Logo
When tackling the RPM and Bongard-Logo problems, we enhance the option pool of a given observed sample by selecting images from two other distinct observed samples. The specific procedure is shown in Figure 10.
@@
Figure 9(a) outlines the approach for RPM problems, and Figure 9(b) illustrates the method for Bongard-Logo problems. Due to computational limitations, we augment the negative examples with three extra samples for RPM and two extra samples for Bongard-Logo.
@@ VII-B NICO
For the NICO dataset, we cannot use external images as negative samples because it would disrupt the data distribution and diminish its significance. This is also why we avoid importing pre-trained model parameters. However, our methodology for improving the quality of negative samples remains unchanged.
When using a basic deep learning model to process NICO data, positive and negative sample comparison is implicitly done through cross-entropy. To enhance the quality of negative samples without extra data or supervision, this paper proposes an explicit comparison method: the Group-based Step-wise Comparison Learning Method (GSCLM). The design of GSCLM is quite straightforward. It separates the training of the perceptual part (convolutional layers) and the inductive part (fully connected layers) of the visual deep learning model, such as Vgg16.
@@ VII-B1 The training of the perceptual part
In the figure, “$\phi$” represents the perceptual part of a visual deep learning network. Subsequently, we impose loss constraints on this representation group to optimize the perceptual component of the deep network. The loss constraints can be expressed as:
@@ VIII Designing More Demanding Negative Examples to Drive Deep-Level Concept Learning (D4C)
We have recognized that the low quality of negative examples hinders deep models from learning precise concepts. To obtain higher-quality negative examples without increasing data or supervision, we propose leveraging the probability feedback provided by the reasoning engine as an evaluation metric to guide the generation of better negative examples. Therefore, a sample generator that produces negative examples with higher evaluations from the reasoning engine is introduced, and its architecture is shown in Figure 12.
Figure 12: The framework that integrates the generator with the reasoning engine
@@
Formula (4) implicitly involves the participation of $K_{n-1}$ and $V_{n-1}$. In this paper, we use the element-wise ELU + 1 function as the activation function $\sigma(\cdot)$. After the step-wise linear attention mechanism calculation is completed, the resulting $O_{s}$ is processed by a feedforward layer, which follows the classic design of the Transformer. This processed output is then summed with $Q^{\prime}_{n}$ to obtain the updated promoter. The updated promoter can either serve as the representation of the generated negative example or act as the input promoter for the block stacked after the current block.
@@
During the training of the generator, it is optimized alternately with the target reasoning engine, with the expectation that the generated negative examples score higher in the engine than the original negative examples. This paper employs the cross-entropy loss function to achieve this learning objective of the generator, enabling it to provide tighter concept boundaries for the reasoning engine, thereby achieving a performance breakthrough for the already trained reasoning engine. Based on the principle that unsupervised learning does not ensure disentanglement [45], the design of such a loss function indicates that this paper is not concerned with the specific attributes or concepts conveyed by the generated negative examples.